OpenStudy (anonymous):

If x = 4 in this equation, what does y equal? 5x + 8y = 4

OpenStudy (anonymous):

5(4)+8y=4 Now just solve for y

jimthompson5910 (jim_thompson5910):

5x + 8y = 4 5(4) + 8y = 4 ... replace every copy of x with 4 20 + 8y = 4 ... ... ... y = ???

OpenStudy (anonymous):

is the answer -2 ?